OpenStudy (anonymous):

The graph of a system of equations with the same slope and the same y-intercepts will have no solutions. a-always b-sometimes c-never

OpenStudy (anonymous):

a solution to a system is where they intersect

OpenStudy (anonymous):

Okay.. I see now.

OpenStudy (anonymous):

y = x + 2 y = x +2

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

this means that there is an infinite amount of solutions

OpenStudy (anonymous):

Thank You.